the sum of twice a number, and 7, is equal to the sum of a number and 6
Answer by ptaylor(2048)

Let x=the number, then we are told that:
2x+7=x+6
2x-x=6-7
x=-1 the number
CK
2*(-1)+7=-1+6
-2+7=-1+6
5=5
Hope this helps---ptaylor